OpenStudy (anonymous):

A squared plus B squared equals C squared. On the triangle one side equals 8 and the other side equals 10. What does the third side equal?

OpenStudy (anonymous):

is the 10 on the longest side?

OpenStudy (anonymous):

yupp

OpenStudy (anonymous):

so c=10, a=8 10^2 = 8^2 + b^2 100 = 64 + b^2 36 = b^2 6 = b (it's a hidden 3,4,5 triangle with a factor of 2)